OpenStudy (mariamikayla):

What is the remainder of dividing the polynomial x^100+x^50-2x^4-x^3+x+1 to the polynomial x^3+x

OpenStudy (mariamikayla):

\[x^{100}+x^{50}-2x^4-x^3+x+1 \ \ \ and \ \ \ x^3+x\]

jhonyy9 (jhonyy9):

x^100 +x^50 -2x^4 -3 +x +1 / x^3 +x = x^97 -x^95 x^100 +x^98 --------------- 0 - x^98 + x^50 - x^98 +x^96 ----------------- 0 - x^96 +x^50 do you can continue it in this way ?

OpenStudy (mariamikayla):

Can i show you afterwards to check my answer,please?

jhonyy9 (jhonyy9):

ok.

OpenStudy (mariamikayla):

Thank you a lot :D

jhonyy9 (jhonyy9):

np.

OpenStudy (phi):

you could use the idea that dividing by (x^3+x) means that the "starting polynomial" P(x) can be written as P(x) = (x^3+x) * Q(x) + R(x) we don't care about Q(x), but we need to find R(x) because R(x) is a remainder it is "smaller" i.e. at least 1 order less than the divisor x^3+x in this case , R(x) is a x^2 + b x + c (an order 2 polynomial) we need to find the coefficients a, b, and c

OpenStudy (phi):

we have P(x) = x^100+x^50-2x^4-x^3+x+1 and P(x) = x(x^2+1) * Q(x) + a x^2 + b x + c next, notice if we pick x=0 P(0) = 0 * Q(0) + a * 0 + b*0 + c or P(0) = c after simplifying. so evaluate P(x) at x=0. we get 1 that means c is 1 P(x) = x(x^2+1) * Q(x) + a x^2 + b x + 1 next, if we pick x= i (this is a complex number , the square root of -1) i^2 + 1 is 0, and we get P(i) = 0 * Q(i) + a* i^2 + bi + 1 P(i) = -a + b i + 1 or P(i) = (1-a) + b i putting in x= i into P(x) we get i^100+i^50-2 i^4- i^3+i+1 I assume you know about complex numbers. i^2 = -1 i^3 = -i and i^4 = 1 using those, we get 1 - 1 - 2 - (-i) + i + 1 which simplifies to -1 + 2i P(i) = -1 + 2 i match that with P(i) = (1-a) + b i b i matches 2 i which means b=2 and 1-a = -1 which means a=2 we finally get R(x) = 2x^2 +2x + 1

OpenStudy (acespeedfighter):

3